OpenStudy (anonymous):

four times the sum of a number and 1 is the same as 5 increased by 5 times the number... find the number.

OpenStudy (anonymous):

4(x+1) = 5+5x ==> x=1

OpenStudy (anonymous):

Because and is another for + as well as increased. IF that helps at all. :)

OpenStudy (anonymous):

so is the answer -9?

OpenStudy (anonymous):

n = -1 4*2 is not equal to 5+5

OpenStudy (anonymous):

can u solve it step by step

OpenStudy (anonymous):

4(x+1)=5+5x 4x+4=5+5x -1=x or x=-1

OpenStudy (anonymous):

ohh thank u so very much i was messing up the sign. :D

OpenStudy (anonymous):

4x (n+1) = 5n+5 4n+4= 5n+5 -1 = n